OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 45 when x = 9. What is value of y when x = 2? y =

OpenStudy (danjs):

y = k*x and (x,y) = (9,45)

OpenStudy (danjs):

find k

OpenStudy (anonymous):

wante sue aleis dmal

OpenStudy (anonymous):

y = k*x 45 = k*9 k = 45/9 k = ? y = kx Put x = 2 and k

OpenStudy (triciaal):

y varies directly see above by @DanJS where k is a constant

OpenStudy (anonymous):

y = 5 mate

OpenStudy (anonymous):

lol thx

OpenStudy (triciaal):

What is value of y when x = 2? y = kx you found k and you have x substitute and compute y

OpenStudy (triciaal):

@Gulivr no that is k

OpenStudy (anonymous):

sorry just kidding y = 10 :P

OpenStudy (anonymous):

?? confused

OpenStudy (anonymous):

see x=9 =1 and y=45 =5 and if x =2 and 2 =1 then y = 10 = 5 :OOO

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

you get me? i simplify

OpenStudy (anonymous):

right